OpenStudy (anonymous):

heLlo, can someone help me out on the following equation. change it to 1st order homogeneous then use v= y/x substitution method xy^2y'=X^3 + Y^3

OpenStudy (anonymous):

i divided both side with xy^2 then the ODE becomes 1st order homogeneous y'= x^3/(xy^2) + y^3/(xy^2) after i worked it out my answer is y = x/(ln |x| +C) but the answer is wrong when compare to the solution manual

OpenStudy (anonymous):

\[xy^2y'=x^3+y^3\] Letting \(y=xv\) (the same substitution, pretty much), you have \(y'=v+xv'\). \[\begin{align*}x^3v^2(v+xv')&=x^3+x^3v^3\\ v^2(v+xv')&=1+v^3\\ v+xv'&=\frac{1+v^3}{v^2}\\ v+xv'&=\frac{1}{v^2}+v\\ xv'&=\frac{1}{v^2}\\ v^2~dv&=\frac{dx}{x} \end{align*}\] Is this the answer you're given (or something similar)? \[y=x\left(\ln (Cx^3)\right)^{1/3}\]